# A module-theoretic interpretation of quantum expansion formula

### v1

@@ 6. THE SKEIN ALGEBRA
A link is a transverse multicurve together with a choice of ordering of the strands at each interior crossing, which means a choice of which strand is ”over” and which is ”under”. Let $K_{q}=K[q^{\pm\frac{1}{2}}]$. Let $K_{q}^{Links}(S)$ denote the free $K_{q}$-module with basis given by the homotopy class $[C]$ of links $C$ in $S$. Then skein module $SK_{q}(S)$ is defined to be $K_{q}^{Links}(S)/R$, where $R$ denote the module of relations in $K_{q}^{Links}(S)$ generated by the following “skein” relation:
@@
The skein module $SK_{q}(S)$ can be an associative algebra using the superposition product:if $X$ and $Y$ are links such that $X\cup Y$ has transverse crossings, then $X\cdot Y$ is equal to $q^{\frac{k}{2}}$ times the link $X\cup Y$ in which strands of $X$ always cross over strands of $Y$ at each crossing. The integer $k$ is defined as follows: for each arc $i$, let $\partial_{1}(i)$ and $\partial_{2}(i)$ denote the two ends of ii(for arbitrary numbering). Then given two arcs $i$, $j$, define
@@
Then $k=\sum\Lambda(i,j)$, where the sum if over all pairs of arcs $i\in X$ and $j\in Y$.
@@
Let $(S,M)$ be an unpunctured surface and $\Gamma$ be a triangulation. Then there is a corresponding skew-symmetric matrix $\Lambda$ satisfying $\Lambda B(\Gamma)=-I$. And we have inclusions
@@
Furthermore, these inclusion identify simlple arcs with cluster variables, boundary arcs being identified with the frozen variables. Triangulations correspond with cluster and mutations of seeds correspond to flips of triangulations. If each component of $S$ contains at least two markings, then the inclusion is actually isomorphism
@@ 7. A MULTIPLICATION FORMULA WITH ONE DIMENSIONAL EXTENSION SPACE
In this section, we provide a multiplication formula between two string modules with one dimensional extension space. We explicitly give all the strings of modules. The extension space was studied in [19].
@@
      if $m$ is a trivial string, then $ac\in I$ and $bd\in I$(whenever they exists, subject to the constraints above).
@@
Let $A$ be a gentle algebra, and $v$ and $w$ be strings. The collection of arrow and overlap extensions between $M(w)$ and $M(v)$ form a basis of $Ext^{1}_{A}(M(v),M(w))$.
@@
    If $v$ is direct, then $v$ is the trivial string corresponding to $e(v)$. Otherwise, $_{c}v$ is obtained from $v$ by deleting the first inverse arrow in $v$ and the direct string preceding it.
@@
    If $v$ is direct, the $v_{h}$ is the trivial string corresponding to $s(v)$. Otherwise, $v_{h}$ is obtained from $v$ by deleting the last inverse arrow in $v$ and the direct string succeeding it.
@@ 7.1. Arrow extension
Let $v$ and $w$ have an arrow extension. Figure 9 shows the arrow extension in surface[21]. In particular, $u_{1}=wa^{-1}v$ is the string of the middle term of extensions. $u_{2}$ is the trivial string corresponding to some vertex $i_{0}$. $u_{3}={{}_{c}}v$ and $u_{4}=v_{h}$.
@@
Then if we identify each arc and the corresponding string module, then $v]=X_{M(v)}$ where we use the same letter for the arc and string. So we have the following:
@@
For any two strings $v$ and $w$ have only one arrow extension. The multiplication formula
@@ 7.2. Overlap extension
Let $v$ and $w$ have an overlap extension. Recall in this case, $v=v_{L}bma^{-1}v_{R}$ and $w=w_{L}d^{-1}mcw_{R}$. Figure 10 shows the overlap extension in surface. From the diagram, we can see $u_{1}=v_{L}bmcw_{R}$ and $u_{2}=w_{L}d^{-1}ma^{-1}v_{R}$.
@@
Then if we identify each arc and the corresponding string module, the following is true:
@@
For any $v$ and $w$ have only one overlap extension. The multiplication formula
@@ 8. The Kronecker Case
For the case when the quiver is $1\rightrightarrows 2$, [20] and [11] both gave the expansion formula of cluster variables in this quantum cluster algebra. In this section, we show they are actually the same.
@@ 8.1. Results in [20]
First, observe that the snake graph corresponding to arcs of this type has alternating face weights(the diagonal edge) 1 and 2.
@@
The snake graph $\mathcal{G}_{n}$ is the following straight snake graph consisting of $2s+1$ tiles of alternating face weight 1 and 2 and with exactly $s+1$ tiles of weight 1 and exactly $s$ tiles of weight 2. The snake graph $\mathcal{H}_{s}$ is obtained from $\mathcal{G}_{s}$ by removing the last tile(with weight 1).
Denote the $2s+1$ tiles of snake graph $\mathcal{G}_{s}$ by
@@
Let $s\geq 0$, define a function on the tiles of snake graph:
@@
Then for snake graph $\mathcal{G}$ which may be either $\mathcal{G}_{s}$ or $\mathcal{H}_{s}$, let $P_{-}(\mathcal{G})$ be the minimal perfect matching. For every perfect matching $P$, the symmetric difference enclosed a set of tiles of $\mathcal{G}$. Let $Twist(P)$ be the set of tiles enclosed. Define a map
@@
These elements play a key role even though they are not cluster variable.
@@ 8.2. Equality
We want to show the expansion formula above equals the expansion formula in [11], which is
@@
Recall that for $P\in\mathcal{P}(\mathcal{G}_{s})$, the symmetric difference actually give the the exponential of $X(P)$, which is the dimension vector $e$ of canonical submodule $M(P)$:
@@
where $g$ is the exponential of $X(P_{-})$. Therefore, if we can show the following equality is true, then the two expansion formulas are equal.
@@
Actually, $u$ and $w$ count the number of tiles with diagonal 1 and 2 in $Twist(P)$. We prove ()(*) by induction on the number of tiles. The equality is obviously true for $\mathcal{G}_{1}$ or $\mathcal{H}_{1}$.
@@ 8.3. In $\mathcal{G}_{s}$
Let $P$ be a perfect matching of $\mathcal{G}_{s}$, if $Twist(P)$ does not contain the last tile, then it can be seen as a perfect matching of $\mathcal{H}_{s}$; if $Twist(P)$ contains the last tile, then it can be seen as a perfect matching of $\mathcal{G}_{s-1}$. The snake graphs and the values of $\alpha$ are showed in Figure 11. They are the snake graph of $\mathcal{G}_{s},\mathcal{H}_{s},\mathcal{G}_{s-1}$ in order, the edges of the minimal perfect matching are showed in blue and the values of $\alpha$ are on the edges.
@@
The correspondence is obvious by definition. We know $P$ and $P^{\prime}$ have the same edge except the last tile label 1. So the dimension vector of $P^{\prime}$ is $(u,w)$. The twist tiles of $P$ and $P^{\prime}$ are the same. Also there are $u$ tiles labeled 1 and $w$ tiles label 2. The values of tiles labeled 1 differ by 1-1 and the values of tiles labeled 2 are the same. Therefore $\alpha(P)=\alpha(P^{\prime})-u$.
For the latter, let $P_{-}$ and $P^{\prime}_{-}$ be the minimal perfect matching of $\mathcal{G}_{s}$ and $\mathcal{H}_{s}$. Then $v(P)-v(P_{-})$ and $v(P^{\prime})-v(P^{\prime}_{-})$ have the same procedure: twist the tile labeled 1 $u$ times and labeled 2 $w$ times. When twisting at tiles labeled 1, they differ by 1-1 and when twisting at tiles labeled 2, they are the same. So $v(P)=v(P^{\prime})-u$
@@
The correspondence is obvious by definition. We know $P$ and $P^{\prime\prime}$ have the same edge except the last two tiles. So the dimension vector of $P^{\prime\prime}$ is (u1,w1)(u-1,w-1). The twist tiles of $P$ and $P^{\prime}$ differ by the last two tiles and $\alpha$ maps these tiles to (s1)+s-(s-1)+s. Also there are u1u-1 tiles labeled 1 and w1w-1 tiles label 2. The values of tiles labeled 1 differ by 1-1 and the values of tiles labeled 2 differ by 1. Therefore $\alpha(P)=\alpha(P^{\prime})+1-(u-1)+w-1=\alpha(P^{\prime})-u+w+1$.
For the latter, let $P_{-}$ and $P^{\prime\prime}_{-}$ be the minimal perfect matching of $\mathcal{G}_{s}$ and $\mathcal{G}_{s-1}$. Then let $Q$ be the perfect matching obtained from $P_{-}$ by twisting at the last two tiles. By computation, $v(Q)=1$. Then v(P)v(Q)v(P)-v(Q) and $v(P^{\prime\prime})-v(P^{\prime\prime}_{-})$ have the same procedure except the last two tiles: twist the tile labeled 1 u1u-1 times and labeled 2 w1w-1 times. When twisting at tiles labeled 1, they differ by 1-1 and when twisting at tiles labeled 2, they differ by $1$. So $v(P)-1=v(P^{\prime\prime})-(u-1)+w-1$. ∎
Proof of (*): Induction on the number of tiles. Assume $\mathcal{H}_{s}$ and $\mathcal{G}_{s-1}$ satisfy the above.
@@ 8.4. In $\mathcal{H}_{s}$
Let $P$ be a perfect matching of $\mathcal{H}_{s}$, if $Twist(P)$ contains the last tile, then it can be seen as a perfect matching of $\mathcal{G}_{s-1}$; if $Twist(P)$ does not contain the last tile, then it can be seen as a perfect matching of $\mathcal{H}_{s-1}$. The snake graph and the values of $\alpha$ are showed in Figure 12. They are the snake graph of $\mathcal{H}_{s},\mathcal{G}_{s-1},\mathcal{H}_{s-1}$ in order, the edges of the minimal perfect matching are showed in blue and the values of $\alpha$ are on the edges.
@@
The correspondence is obvious by definition. We know $P$ and $P^{\prime}$ have the same edge except the last tile label 1. So the dimension vector of $P^{\prime}$ is (u,w1)(u,w-1). The twist tiles of $P$ and $P^{\prime}$ differ by the last tiles and $\alpha$ maps this tile to s+1-s+1. Also there are $u$ tiles labeled 1 and w1w-1 tiles label 2. The values of tiles labeled 1 are the same and the values of tiles labeled 2 differ by $1$. Therefore $\alpha(P)=\alpha(P^{\prime})-s+1+w-1$.
For the latter, let $P_{-}$ and $P^{\prime}_{-}$ be the minimal perfect matching of $\mathcal{H}_{s}$ and $\mathcal{G}_{s-1}$. Then let $Q$ be the perfect matching obtained from $P_{-}$ by twisting at the last tile. By computation, v(Q)=s1v(Q)=s-1. Then v(P)v(Q)v(P)-v(Q) and $v(P^{\prime})-v(P^{\prime}_{-})$ have the same procedure: twist the tile labeled 1 $u$ times and labeled 2 w1w-1 times. When twisting at tiles labeled 1, they are the same and when twisting at tiles labeled 2, they differ by 1-1. So $v(P)=v(P^{\prime})+s-1-(w-1)$
@@
The correspondence is obvious by definition. We know $P$ and $P^{\prime\prime}$ have the same edge except the last two tiles. So the dimension vector of $P^{\prime\prime}$ is $(u,w)$. The twist tiles of $P$ and $P^{\prime\prime}$ are the same and there are $u$ tiles labeled 1 and $w$ tiles label 2. The values of tiles labeled 1 differ by 1-1 and the values of tiles labeled 2 differ by 1. Therefore $\alpha(P)=\alpha(P^{\prime})-u+w$.
For the latter, let $P_{-}$ and $P^{\prime\prime}_{-}$ be the minimal perfect matching of $\mathcal{G}_{s}$ and $\mathcal{G}_{s-1}$. Then $v(P)-v(P_{-})$ and $v(P^{\prime\prime})-v(P^{\prime\prime}_{-})$ have the same procedure: twist the tile labeled 1 $u$ times and labeled 2 $w$ times. When twisting at tiles labeled 1, they differ by $1$ and when twisting at tiles labeled 2, they differ by 1-1. So $v(P)-1=v(P^{\prime\prime})+u-w$. ∎
Proof of (*): Induction on the number of tiles. Assume $\mathcal{G}_{s-1}$ and $\mathcal{H}_{s-1}$ satisfy the above.

### v2

@@ 6. THE SKEIN ALGEBRA
A link is a transverse multicurve together with a choice of ordering of the strands at each interior crossing, which means a choice of which strand is “over” and which is “under”. Let $K_{q}=K[q^{\pm\frac{1}{2}}]$. Let $K_{q}^{Links}(S)$ denote the free $K_{q}$-module with basis given by the homotopy classes $[C]$ of links $C$ in $S$. Then the skein module $SK_{q}(S)$ is defined to be $K_{q}^{Links}(S)/R$, where $R$ denotes the module of relations in $K_{q}^{Links}(S)$ generated by the following “skein” relation:
@@
The skein module $SK_{q}(S)$ is an associative algebra using the superposition product: if $X$ and $Y$ are links such that $X\cup Y$ has transverse crossings, then $X\cdot Y$ is equal to $q^{\frac{k}{2}}$ times the link $X\cup Y$ in which strands of $X$ always cross over strands of $Y$ at each crossing. The integer $k$ is defined as follows: for each arc $i$, let $\partial_{1}(i)$ and $\partial_{2}(i)$ denote the two ends of $i$ (for arbitrary numbering). Then given two arcs $i$, $j$, define
@@
Then $k=\sum\Lambda(i,j)$, where the sum is over all pairs of arcs $i\in X$ and $j\in Y$.
@@
Let $(S,M)$ be an unpunctured surface and $\Gamma$ be a triangulation. Then there is a corresponding skew-symmetric matrix $\Lambda$ satisfying $\Lambda B(\Gamma)=-I$. We have an inclusion
@@
Furthermore, this inclusion identifies simple arcs with cluster variables, boundary arcs being identified with the frozen variables. Triangulations correspond with clusters and mutations of seeds correspond to flips of triangulations. If each component of $S$ contains at least two markings, then the inclusion is actually an isomorphism.
@@ 7. A MULTIPLICATION FORMULA WITH ONE DIMENSIONAL EXTENSION SPACE
In this section, we provide a multiplication formula between two string modules with one-dimensional extension space. We explicitly give all the strings of these modules. The extension space was studied in [12].
@@
      if $m$ is a trivial string, then $ac\in I$ and $bd\in I$ (whenever they exist, subject to the constraints above).
@@
Let $A$ be a gentle algebra, and $v$ and $w$ be strings. The collection of arrow and overlap extensions between $M(w)$ and $M(v)$ form a basis for $Ext^{1}_{A}(M(v),M(w))$.
@@
    If $v$ is direct, then $_{c}v$ is the trivial string corresponding to $e(v)$. Otherwise, $_{c}v$ is obtained from $v$ by deleting the first inverse arrow in $v$ and the direct string preceding it.
@@
    If $v$ is direct, then $v_{h}$ is the trivial string corresponding to $s(v)$. Otherwise, $v_{h}$ is obtained from $v$ by deleting the last inverse arrow in $v$ and the direct string succeeding it.
@@ 7.1. Arrow extension
Let $v$ and $w$ have an arrow extension. Figure 9 shows the arrow extension on the surface [9]. In particular, $u_{1}=wa^{-1}v$ is the string of the middle term of the extension. $u_{2}$ is the trivial string corresponding to some vertex $i_{0}$. $u_{3}={{}_{c}}v$ and $u_{4}=v_{h}$.
@@
If we identify each arc with the corresponding string module, then $v]=X_{M(v)}$ where we use the same letter for the arc and string. So we have the following:
@@
If two strings $v$ and $w$ have only one arrow extension, then we have a multiplication formula
@@ 7.2. Overlap extension
Let $v$ and $w$ have an overlap extension. Recall that in this case, $v=v_{L}bma^{-1}v_{R}$ and $w=w_{L}d^{-1}mcw_{R}$. Figure 10 shows the overlap extension on the surface. From the diagram, we can see $u_{1}=v_{L}bmcw_{R}$ and $u_{2}=w_{L}d^{-1}ma^{-1}v_{R}$.
@@
Then if we identify each arc with the corresponding string module, the following is true:
@@
If two strings $v$ and $w$ have only one overlap extension, then we have a multiplication formula
@@ 8. The Kronecker Case
For the case when the quiver is $1\rightrightarrows 2$, [10] and [18] both gave the expansion formula of cluster variables in this quantum cluster algebra. In this section, we show that they are actually the same.
@@ 8.1. Results in [10]
First, observe that the snake graph corresponding to arcs of this type has alternating face weights (diagonal labels) 1 and 2.
@@
The snake graph $\mathcal{G}_{n}$ is the following straight snake graph consisting of $2s+1$ tiles with alternating face weights $1$ and $2$, with exactly $s+1$ tiles of weight $1$ and exactly $s$ tiles of weight $2$. The snake graph $\mathcal{H}_{s}$ is obtained from $\mathcal{G}_{s}$ by removing the last tile (with weight 1).
Denote the $2s+1$ tiles of the snake graph $\mathcal{G}_{s}$ by
@@
Let $s\geq 0$, define a function on the tiles of the snake graph:
@@
Then for snake graph $\mathcal{G}$ which may be either $\mathcal{G}_{s}$ or $\mathcal{H}_{s}$, let $P_{-}(\mathcal{G})$ be the minimal perfect matching. For every perfect matching $P$, the symmetric difference encloses a set of tiles of $\mathcal{G}$. Let $Twist(P)$ be the set of tiles enclosed. Define a map
@@
These elements play a key role even though they are not cluster variables.
@@ 8.2. Equality
We want to show that the expansion formula above equals the expansion formula in [18], which is
@@
Recall that for $P\in\mathcal{P}(\mathcal{G}_{s})$, the symmetric difference actually gives the exponent of $X(P)$, which is the dimension vector $e$ of canonical submodule $M(P)$:
@@
where $g$ is the exponent of $X(P_{-})$. Therefore, if we can show the following equality is true, then the two expansion formulas are equal.
@@
Actually, $u$ and $w$ count the number of tiles with diagonal 1 and 2 in $Twist(P)$. We prove ()(*) by induction on the number of tiles. The equality is obviously true for $\mathcal{G}_{1}$ and $\mathcal{H}_{1}$.
@@ 8.3. In $\mathcal{G}_{s}$
Let $P$ be a perfect matching of $\mathcal{G}_{s}$. If $Twist(P)$ does not contain the last tile, then it can be viewed as a perfect matching of $\mathcal{H}_{s}$; if $Twist(P)$ contains the last tile, then it can be viewed as a perfect matching of $\mathcal{G}_{s-1}$. The snake graphs and the values of $\alpha$ are shown in Figure 11. They are the snake graph of $\mathcal{G}_{s},\mathcal{H}_{s},\mathcal{G}_{s-1}$ in order, the edges of the minimal perfect matching are shown in blue and the values of $\alpha$ are on the edges.
@@
The correspondence is obvious by definition. We know $P$ and $P^{\prime}$ have the same edge except for the last tile labeled 1. So the dimension vector of $P^{\prime}$ is $(u,w)$. The twist tiles of $P$ and $P^{\prime}$ are the same. Also, there are $u$ tiles labeled 1 and $w$ tiles labeled 2. The values of tiles labeled 1 differ by 1-1 and the values of tiles labeled 2 are the same. Therefore $\alpha(P)=\alpha(P^{\prime})-u$.
For the latter, let $P_{-}$ and $P^{\prime}_{-}$ be the minimal perfect matchings of $\mathcal{G}_{s}$ and $\mathcal{H}_{s}$. Then $v(P)-v(P_{-})$ and $v(P^{\prime})-v(P^{\prime}_{-})$ have the same procedure: twist the tile labeled 1 $u$ times and labeled 2 $w$ times. When twisting at tiles labeled 1, they differ by 1-1 and when twisting at tiles labeled 2, they are the same. So $v(P)=v(P^{\prime})-u$. ∎
@@
The correspondence is obvious by definition. We know $P$ and $P^{\prime\prime}$ have the same edges except for the last two tiles. So the dimension vector of $P^{\prime\prime}$ is (u1,w1)(u-1,w-1). The twist tiles of $P$ and $P^{\prime\prime}$ differ by the last two tiles and $\alpha$ maps these tiles to (s1)+s-(s-1)+s. Also there are u1u-1 tiles labeled 1 and w1w-1 tiles labeled 2. The values of tiles labeled 1 differ by 1-1 and the values of tiles labeled 2 differ by 1. Therefore $\alpha(P)=\alpha(P^{\prime\prime})+1-(u-1)+w-1=\alpha(P^{\prime\prime})-u+w+1$.
For the latter, let $P_{-}$ and $P^{\prime\prime}_{-}$ be the minimal perfect matchings of $\mathcal{G}_{s}$ and $\mathcal{G}_{s-1}$. Then let $Q$ be the perfect matching obtained from $P_{-}$ by twisting at the last two tiles. By computation, $v(Q)=1$. Then v(P)v(Q)v(P)-v(Q) and $v(P^{\prime\prime})-v(P^{\prime\prime}_{-})$ have the same procedure except the last two tiles: twist the tile labeled 1 u1u-1 times and labeled 2 w1w-1 times. When twisting at tiles labeled 1, they differ by 1-1 and when twisting at tiles labeled 2, they differ by $1$. So $v(P)-1=v(P^{\prime\prime})-(u-1)+w-1$. ∎
Proof of (*): Induction on the number of tiles. Assume that $\mathcal{H}_{s}$ and $\mathcal{G}_{s-1}$ satisfy the above.
@@ 8.4. In $\mathcal{H}_{s}$
Let $P$ be a perfect matching of $\mathcal{H}_{s}$. If $Twist(P)$ contains the last tile, then it can be seen as a perfect matching of $\mathcal{G}_{s-1}$; if $Twist(P)$ does not contain the last tile, then it can be seen as a perfect matching of $\mathcal{H}_{s-1}$. The snake graph and the values of $\alpha$ are shown in Figure 12. They are the snake graph of $\mathcal{H}_{s},\mathcal{G}_{s-1},\mathcal{H}_{s-1}$ in order, the edges of the minimal perfect matching are shown in blue and the values of $\alpha$ are on the edges.
@@
The correspondence is obvious by definition. We know $P$ and $P^{\prime}$ have the same edges except for the last tile labeled 1. So the dimension vector of $P^{\prime}$ is (u,w1)(u,w-1). The twist tiles of $P$ and $P^{\prime}$ differ by the last tile and $\alpha$ maps this tile to s+1-s+1. Also there are $u$ tiles labeled 1 and w1w-1 tiles labeled 2. The values of tiles labeled 1 are the same and the values of tiles labeled 2 differ by $1$. Therefore $\alpha(P)=\alpha(P^{\prime})-s+1+w-1$.
For the latter, let $P_{-}$ and $P^{\prime}_{-}$ be the minimal perfect matchings of $\mathcal{H}_{s}$ and $\mathcal{G}_{s-1}$. Then let $Q$ be the perfect matching obtained from $P_{-}$ by twisting at the last tile. By computation, v(Q)=s1v(Q)=s-1. Then v(P)v(Q)v(P)-v(Q) and $v(P^{\prime})-v(P^{\prime}_{-})$ have the same procedure: twist the tile labeled 1 $u$ times and labeled 2 w1w-1 times. When twisting at tiles labeled 1, they are the same and when twisting at tiles labeled 2, they differ by 1-1. So $v(P)=v(P^{\prime})+s-1-(w-1)$. ∎
@@
The correspondence is obvious by definition. We know $P$ and $P^{\prime\prime}$ have the same edges except for the last two tiles. So the dimension vector of $P^{\prime\prime}$ is $(u,w)$. The twist tiles of $P$ and $P^{\prime\prime}$ are the same and there are $u$ tiles labeled 1 and $w$ tiles labeled 2. The values of tiles labeled 1 differ by 1-1 and the values of tiles labeled 2 differ by 1. Therefore $\alpha(P)=\alpha(P^{\prime\prime})-u+w$.
For the latter, let $P_{-}$ and $P^{\prime\prime}_{-}$ be the minimal perfect matchings of $\mathcal{G}_{s}$ and $\mathcal{G}_{s-1}$. Then $v(P)-v(P_{-})$ and $v(P^{\prime\prime})-v(P^{\prime\prime}_{-})$ have the same procedure: twist the tile labeled 1 $u$ times and labeled 2 $w$ times. When twisting at tiles labeled 1, they differ by $1$ and when twisting at tiles labeled 2, they differ by 1-1. So $v(P)-1=v(P^{\prime\prime})+u-w$. ∎
Proof of (*): Induction on the number of tiles. Assume that $\mathcal{G}_{s-1}$ and $\mathcal{H}_{s-1}$ satisfy the above.